# Trajectory-Based Multi-Objective Hyperparameter Optimization for Model Retraining

### v1

@@ Abstract
Training machine learning models inherently involves a resource-intensive and noisy iterative learning procedure that allows epoch-wise monitoring of the model performance. However, in multi-objective hyperparameter optimization scenarios, the insights gained from the iterative learning procedure typically remain underutilized. We notice that tracking the model performance across multiple epochs under a hyperparameter setting creates a trajectory in the objective space and that trade-offs along the trajectories are often overlooked despite their potential to offer valuable insights to decision-making for model retraining. Therefore, in this study, we propose to enhance the multi-objective hyperparameter optimization problem by having training epochs as an additional decision variable to incorporate trajectory information. Correspondingly, we present a novel trajectory-based multi-objective Bayesian optimization algorithm characterized by two features: 1) an acquisition function that captures the improvement made by the predictive trajectory of any hyperparameter setting and 2) a multi-objective early stopping mechanism that determines when to terminate the trajectory to maximize epoch efficiency. Numerical experiments on diverse synthetic simulations and hyperparameter tuning benchmarks indicate that our algorithm outperforms the state-of-the-art multi-objective optimizers in both locating better trade-offs and tuning efficiency.
@@ 1 Introduction
With the expanding complexity of machine learning (ML) models, there is a significant surge in the demand for Hyperparameter Optimization (HPO). This surge is not only in pursuit of model prediction accuracy but also for ensuring the computational efficiency and robustness of models in real-world scenarios, which leads to the Multi-Objective Hyperparameter Optimization (MOHPO) [12, 21, 28]. Addressing HPO has long been challenging as it involves resource-intensive model training that prevents optimizers from exhaustively exploring the hyperparameter space. In this context, Bayesian Optimization (BO) has become increasingly popular [4, 36]. This approach builds a probabilistic surrogate model, e.g., Gaussian Process (GP) [33], for the objective and samples a new solution by maximizing an acquisition function formulated by the prediction and uncertainty of the surrogate model. Due to the sample efficiency of BO, many research works have extended it in different ways to Multi-Objective BO (MOBO) and successfully applied to solve MOHPO [1, 8, 9].
Nevertheless, conventional BO methods require observing the model performance that is fully trained after a maximum number of epochs, which could potentially lead to a waste of computational resources if early indications suggest sub-optimal performance. In general, the training behind many ML models is an iterative learning procedure where a gradient-based optimizer updates the model epoch by epoch. This procedure allows users to delineate a learning curve for any hyperparameter setting by epoch-wisely monitoring the intermediate model performance (see Figure 1(A) and (B)), benefiting from which the literature has presented a set of epoch-efficient single-objective BO methods [38, 7, 29, 3] to avoid computational waste. Unfortunately, to the best of our knowledge, there remains a gap in multi-objective research that fully accounts for iterative learning procedures.
@@
Considering MOHPO with iterative learning procedures motivates the critical question of whether a trade-off that achieves the best model performance under a fixed training cost exists when training with fewer epochs than the maximum allowed. Figure 1(C) depicts such a bi-objective scenario, where we extend the concept of the learning curve by introducing “trajectory” to describe the iterative training procedure. For instance, the trajectory of a $d$-dimensional hyperparameter setting $\bm{x}_{1}$ consists of its model performances (training cost and validation loss) observed from epoch $t=1$ to $10$. The emergence of trajectories reveals the limitation of traditional MOHPO in that it focuses only on fully-trained model performances ($t=10$) and ignores a large amount of partially-trained model performances ($t<10$) that can also contribute to trade-offs. Nevertheless, these trade-offs offer valuable insights for decision-making in scenarios where ML models are periodically retrained on similar datasets with similar performance and need rapid deployment, such as recurrent data analyst jobs in the cloud [5, 27] and data drift detection in self-adaptive systems [25, 6]. To avoid repetitive and costly tuning in these cases, it is a common practice to use the same hyperparameter setting for every retraining [44, 21], which makes it important to understand at the initial stage how many epochs to retrain to strike the desired balance between objectives. Consequently, in this study, we propose to tune the hyperparameter setting together with the number of training epochs and formulate an Enhanced MOHPO (i.e., EMOHPO).
To effectively solve EMOHPO, it is necessary to develop a new MOBO method capable of integrating the information drawn from trajectories into sequential sampling. Given that a trajectory may contribute more than one trade-off (illustrated in Figure 1(C)), assessing the quality of a hyperparameter setting in EMOHPO requires a comprehensive consideration of its entire trajectory. Existing multi-objective acquisition functions tend to favor hyperparameter settings with superior model performance at a single position, which potentially leads to trajectories with limited trade-offs. Furthermore, similar to single-objective scenarios, as an iterative learning procedure proceeds, we can epoch-wisely update predictions for the unobserved segment of the trajectory and decide on termination to improve algorithm efficiency. However, it is crucial to guarantee that such an early stopping mechanism does not compromise the optimization outcomes. In the context of EMOHPO, this implies that only after observing as many trade-offs along the trajectory as possible can the iterative learning procedure be stopped.
@@
Contributions: (1) For the first time we formulate EMOHPO, a variant of MOHPO enhanced by the trajectory information, to account for the evolution of model performance across epochs and hence uncover trade-offs that are often overlooked in conventional MOHPO studies. (2) We introduce a novel trajectory-based MOBO method for EMOHPO. The proposed method samples the next hyperparameter setting based on the quality of its anticipated trajectory and subsequently determines when to terminate the trajectory to maximize epoch efficiency. (3) Through comprehensive experiments on synthetic simulations and machine learning benchmarks, we demonstrate the effectiveness and efficiency of our method in identifying trade-offs while conserving computational resources.
2 Enhanced Multi-Objective Hyperparameter Optimization Problem
@@
Throughout the paper, we consider the sequential minimization of an Enhanced Multi-Objective Hyperparameter Optimization problem (EMOHPO) in the following form:
@@
$\min_{(\bm{x},t)\in\mathbb{X}\times\mathbb{T}}\bm{f}(\bm{x},t)=\left[f_{1}(\bm{x},t),\dots,f_{k}(\bm{x},t)\right],$ (1)
where $\bm{f}$ comprises $k$ objective functions, each of which represents a different performance measure of an ML model, e.g., validation loss and training cost. Each $\bm{x}$ is a $d$-dimensional hyperparameter setting within a compact set $\mathbb{X}\subset\mathbb{R}^{d}$. We further consider an iterative learning procedure, where, with any feasible hyperparameter setting $\bm{x}$, we can train the ML model for any epoch $t\in\mathbb{T}=\{1,\dots,t_{max}\}$ and observe the intermediate but noisy performance $\bm{y}(\bm{x},t)=\left[y_{1}(\bm{x},t),\dots,y_{k}(\bm{x},k)\right]$ where $y_{i}(\bm{x},t)=f_{i}(\bm{x},t)+\varepsilon_{i}$ and $\varepsilon_{i}\sim\mathcal{N}(0,\sigma_{i}^{2})$ with variance $\sigma_{i}^{2}$ for any $i\in\{1,\dots,k\}$. Let $\bm{z}=(\bm{x},t)$ denote any feasible query pair and $\mathbb{Z}=\mathbb{X}\times\mathbb{T}$ the overall feasible domain. As per Definitions 1 and 2 below, solving the EMOHPO in (1) is equivalent to locating a set of trade-offs that approximate the Pareto-optimal set and front as accurately as possible. When making the final decision, for example, among the trade-offs of EMOHPO shown in Figure 1(C), one can easily select the best pair $(\bm{x},t)$ referring to training the ML model with hyperparameter setting $\bm{x}$ for $t$ epochs - to achieve the desired level of validation loss within a feasible time budget.
@@
The Pareto-optimal set $Z^{*}$ is composed of query pair $\bm{z}$ whose $\bm{f}(\bm{z})$ is not dominated by $\bm{f}(\bm{z}^{\prime})$ of any other $\bm{z^{\prime}}\in\mathbb{Z}$, i.e., $Z^{*}=\{\bm{z}\mid\bm{z}\in\mathbb{Z},\nexists\bm{z}^{\prime}\in\mathbb{Z},\bm{f}(\bm{z}^{\prime})\prec\bm{f}(\bm{z})\}$. The corresponding set of solutions $F^{*}=\{\bm{f}(\bm{z})\mid\bm{z}\in Z^{*}\}$ is referred to as Pareto-optimal front.
In the context of iterative learning, each objective function $f_{i}(\bm{x},\cdot)$, for $i\in\{1,\dots,k\}$, can be treated as a learning curve associated with hyperparameter setting $\bm{x}$. However, a learning curve only enables the analysis of one performance measure at a time. To holistically analyze multiple performance measures, we introduce the concept of trajectory and define it as the collection of all observable model performances during the training with hyperparameter setting $\bm{x}$, i.e., $Trj(\bm{x}):=\{\bm{f}(\bm{x},t)\}_{t=1}^{t_{max}}=\{\left[f_{1}(\bm{x},t),\dots f_{k}(\bm{x},t)\right]\}_{t=1}^{t_{max}}$. Inherently, a trajectory encapsulates the information provided by multiple learning curves.
As demonstrated in Section 1, EMOHPO is an enhancement of the standard MOHPO with the information provided by trajectories. More specifically, compared to (1), MOHPO restricts its identification of trade-offs among the fully-trained model performances and is expressed as,
$\min_{\bm{x}\in\mathbb{X}}\bm{f}(\bm{x},t_{max})=\left[f_{1}(\bm{x},t_{max}),\dots,f_{k}(\bm{x},t_{max})\right],$ (2)
where the second input $t_{max}$ is a fixed constant and is always assumed to be implicit in earlier studies’ formulations [23, 8, 9]. Taking the second input as an explicit variable $t$ significantly distinguishes EMOHPO from standard MOHPO in two aspects: (I) EMOHPO allows an optimizer to collect observations from $\bm{f}(\bm{x},1)$ to $\bm{f}(\bm{x},t)$ when querying at $(\bm{x},t)$ because the value of $\bm{f}(\bm{x},t^{\prime})$ is observed before $\bm{f}(\bm{x},t)$ for any $t^{\prime}<t$. This feature necessitates the design of a more sample-efficient strategy, which aims to avoid repeatedly visiting the same hyperparameter setting across different epoch numbers. (II) EMOHPO broadens the optimization domain from $\mathbb{X}\times\{t_{max}\}$ to $\mathbb{X}\times\mathbb{T}$ and accounts for all observations on the trajectories. As a consequence, the Pareto-optimal front of EMOHPO is superior to or at least equivalent to that of MOHPO because the former additionally captures trade-offs that may emerge during iterative learning (see Figure 1(C)).
@@ 3 Related Work in Bayesian Optimization
Bayesian Optimization for Iterative Learning: By appropriately characterizing the learning curve, epoch-efficient BO aims to predict the fully-trained model performance based on a partially observed learning curve to avoid ineffective epochs of training. Freeze-Thaw BO [38] introduces GP with an exponential decaying kernel to model the validation loss over time and allows the training to be paused and later resumed under hyperparameter settings that show promise. BOHB [14], a BO extension of HyperBand [24], allocates training epochs through random sampling and utilizes successive halving to eliminate suboptimal hyperparameter settings. Instead of dynamically allocating computational budget among hyperparameter settings, BO-BOS [7] combines BO with Bayesian optimal stopping to early stop the training with a hyperparameter setting predicted to yield poor model performance. Similarly, both BOIL [29] and BAPI [3] incorporate strategies for early stopping with a particular focus on considering the learning curve of the training cost. BOIL integrates the cost into the acquisition function to prioritize cost-effective ML training procedures, whereas BAPI imposes a fixed total budget over cost. Unfortunately, there has been no epoch-efficient BO designed for MOHPO or EMOHPO.
@@
Multi-Objective Bayesian Optimization: Many MOBO methods have been developed for multi-objective HPO by extending the vanilla BO framework. These methods generally build a surrogate model for each objective to minimize the necessity of actual resource-intensive objective evaluations and they differ in the implementation of acquisition function. A straightforward approach is to convert a multiple-objective problem into a single-objective problem through techniques like random scalarization, which allows a direct application of standard acquisition functions. For example, ParEGO [23] and TS-TCH [30] respectively apply Expected Improvement (EI) [18] and Thompson Sampling (TS) [40]. However, this approach often encounters limitations in adequately exploring the Pareto-optimal front. Therefore, acquisition functions biased towards the Pareto-optimal front, such as Expected Hypervolume Improvement (EHVI) [13], have been designed. Despite the effectiveness and popularity of EHVI, its computational intensity presents a significant challenge in the development of BO methods [17]. More recently, $q$EHVI [8] and $q$NEHVI [9] have extended EHVI for parallel multi-point selection and batch optimization through Monte Carlo (MC) integration [13] and they have demonstrated notable empirical performance. Alternatives to EHVI can be found in [16, 1, 37, 45].
@@
Multi-Fidelity Bayesian Optimization: Multi-fidelity BO has a rich research history [19, 20, 35, 39] and it facilitates the optimization of fully-trained model performance by utilizing its low-fidelity approximations which can be obtained either by using a partial training dataset or by limiting the number of training epochs. Although multi-fidelity BO shares similarities with epoch-efficient BO, it predetermines the fidelity level before initiating the iterative learning procedure and therefore ignores the observations during the procedure. For example, FABOLAS [22] considers the data subset size as the fidelity level and jointly selects the hyperparameter setting and the data subset for model training. BOCA [20] expands the discrete fidelity space to continuous for a more general problem setting. Multi-fidelity MOBO has also been studied in the literature [2]; however, it cannot be applied to solving EMOHPO because the fidelity level is no longer an additional degree of freedom in (4), which makes low-fidelity approximations not available.

### v2

# A Trajectory-Based Bayesian Approach to Multi-Objective Hyperparameter Optimization with Epoch-Aware Trade-Offs

@@ Abstract
Training machine learning models inherently involves a resource-intensive and noisy iterative learning procedure that allows epoch-wise monitoring of the model performance. However, the insights gained from the iterative learning procedure typically remain underutilized in multi-objective hyperparameter optimization scenarios. Despite the limited research in this area, existing methods commonly identify the trade-offs only at the end of model training, overlooking the fact that trade-offs can emerge at earlier epochs in cases such as overfitting. To bridge this gap, we propose an enhanced multi-objective hyperparameter optimization problem that treats the number of training epochs as a decision variable, rather than merely an auxiliary parameter, to account for trade-offs at an earlier training stage. To solve this problem and accommodate its iterative learning, we then present a trajectory-based multi-objective Bayesian optimization algorithm characterized by two features: 1) a novel acquisition function that captures the improvement along the predictive trajectory of model performances over epochs for any hyperparameter setting and 2) a multi-objective early stopping mechanism that determines when to terminate the training to maximize epoch efficiency. Experiments on synthetic simulations and hyperparameter tuning benchmarks demonstrate that our algorithm can effectively identify the desirable trade-offs while improving tuning efficiency.
@@ 1 Introduction
With the expanding complexity of machine learning (ML) models, there is a significant surge in the demand for Hyperparameter Optimization (HPO). This surge is not only in pursuit of model prediction accuracy but also for ensuring the computational efficiency and robustness of models in real-world scenarios, which leads to the optimization task of finding the trade-off hyperparameter settings among multiple competing objectives $\bm{f}=\{f_{1},\dots,f_{k}\}$, known as Multi-Objective Hyperparameter Optimization (MOHPO) [Eggensperger et al., 2021, Karl et al., 2023, Morales-Hernández et al., 2023]. While MOHPO focuses solely on tuning hyperparameters $\bm{x}$, we extend this framework at its core by jointly tuning $\bm{x}$ and training epochs $t$, i.e., effectively optimizing objectives $\bm{f}(\bm{x},t)$, to uncover superior trade-offs that emerge during iterative training (see Section 2 for more details).
Addressing HPO has long been challenging as it involves resource-intensive model training that prevents optimizers from exhaustively exploring the hyperparameter space. In this context, Bayesian Optimization (BO) has become increasingly popular [Srinivas et al., 2009, Bergstra et al., 2011, Lévesque et al., 2016, Foldager et al., 2023]. This approach builds a probabilistic surrogate model, e.g., Gaussian Process (GP) [Rasmussen and Williams, 2005], for the objective and samples a new solution by maximizing an acquisition function formulated by the prediction and uncertainty of the surrogate model. Nevertheless, traditional BO methods require observing the model performance that is fully trained after a maximum number of epochs, which could potentially lead to a waste of computational resources if early indications suggest sub-optimal performance. In general, the training behind many ML models is an iterative learning procedure where a gradient-based optimizer updates the model epoch by epoch. This procedure allows users to delineate a learning curve for any hyperparameter setting by epoch-wisely monitoring the intermediate model performance (see Figure 1(A) and (B)), benefiting from which prior research has introduced a set of epoch-efficient single-objective BO methods [Swersky et al., 2014, Dai et al., 2019, Nguyen et al., 2020, Belakaria et al., 2023] to avoid computational waste.
@@
The concept of leveraging iterative learning to achieve more granular control over training epochs can also be extended to MOHPO. Existing methods in multi-fidelity MOHPO [Belakaria et al., 2020, Schmucker et al., 2021] offer a relevant perspective by treating training epoch as an auxiliary parameter and aim to use the information obtained from lower fidelities (i.e., fewer training epochs) to facilitate the search for the trade-offs at the highest fidelity (i.e., the maximum training epoch). The key assumption underlying this stream of research is that higher fidelity levels generally lead to more desirable model performance. However, this assumption does not always hold in the context of ML training. For example, in the case of overfitting, a model trained for fewer epochs can outperform a fully-trained model, making it the desired outcome. This further leads to a critical question: does a trade-off among multiple objectives emerge when the number of training epochs is fewer than the maximum allowed?
Figure 1(C) depicts the objective space of an HPO with two objectives (training cost and validation loss), where each $d$-dimensional hyperparameter setting $\bm{x}$ is trained for up to $10$ epochs, with its fully-trained model performance denoted by a star symbol. To better capture these dynamics, we extend the concept of the learning curve by introducing the notion of a “trajectory” to describe the evolution of model performance across epochs during the iterative training procedure. For instance, the trajectory of $\bm{x}_{1}\in\mathbb{R}^{d}$ (denoted by the pink curve) consists of its model performances observed from epoch $t=1$ to $10$. The emergence of trajectories highlights the limitation of multi-fidelity MOHPO in that their optimality is limited to fully-trained model performances only, ignoring a large amount of partially-trained model performances ($t<10$) that can also contribute to trade-offs.
To address the aforementioned limitation, in this study, we propose to jointly tune the hyperparameter setting and training epoch in the context of iterative learning, and formulate a novel Enhanced MOHPO (i.e., EMOHPO) whose optimization target is to uncover the trade-offs across all partially- and fully-trained model performances, or equivalently, the trade-offs over trajectories (depicted by the shaded points in Figure 1(C)). In fact, one important application of EMOHPO is to avoid the overfitting issue. The trade-offs of EMOHPO provide valuable insights for determining how many epochs should be allocated to achieve better generalization. Additionally, in scenarios where ML models are periodically retrained on similar datasets and need rapid deployment, such as recurrent data analyst jobs in the cloud [Casimiro et al., 2020, Mendes et al., 2020] and data drift detection in self-adaptive systems [Mahadevan and Mathioudakis, 2024, Casimiro et al., 2024], the decision maker can benefit from the trade-offs of EMOHPO to gain a better understanding of the optimal hyperparameter setting and training epoch to strike the desired balance between objectives while avoiding repetitive and costly tuning for each retraining cycle.
Meanwhile, it is important to note that while the training epoch in EMOHPO can be interpreted as a fidelity level, it is explicitly treated as a decision variable. This distinction makes the optimization methods designed for multi-fidelity MOHPO incompatible with EMOHPO, as they seek optimal hyperparameter settings at the highest fidelity level only and do not take training epoch as part of the decision-making process. To this end, we introduce a Trajectory-based MOBO (i.e., TMOBO) algorithm, which is designed to fully leverage the trajectory information for sequential sampling and granular control over training epochs. More specifically, TMOBO samples a hyperparameter setting in each iteration, as is common in BO methods; however, we introduce a specialized acquisition function that accounts for the contribution of the entire trajectory of model performances, instead of any single one of them, associated with a hyperparameter setting. This extension is crucial as each trajectory may contribute to multiple trade-offs (illustrated in Figure 1(C)). Then, during the iterative learning of the sampled hyperparameter setting, TMOBO epoch-wisely updates predictions for the unobserved segment of the trajectory and decides on termination to improve algorithm efficiency. The proposed early stopping mechanism additionally ensures that the iterative learning continues until sufficient trade-offs along the trajectory have been collected, making it significantly different from the stopping criteria used in previous studies.
@@
Contributions: (1) For the first time we formulate EMOHPO to account for the evolution of model performance across epochs and hence uncover trade-offs over trajectories which are often overlooked in multi-fidelity MOHPO studies. (2) We introduce a trajectory-based MOBO method to solve EMOHPO. The proposed method samples the next hyperparameter setting using a novel acquisition function that encapsulates trajectory information and determines when to terminate the iterative learning by a conservative early stopping mechanism. (3) Through comprehensive experiments on synthetic simulations and machine learning benchmarks, we demonstrate the effectiveness and efficiency of our method in identifying trade-offs while conserving computational resources.
2 Enhanced Multi-Objective Hyperparameter Optimization
@@
Throughout the paper, we consider the sequential minimization of an Enhanced Multi-Objective Hyperparameter Optimization problem (EMOHPO) formulated as follows,
@@
$\displaystyle\min_{(\bm{x},t)\in\mathbb{X}\times\mathbb{T}}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \bm{f}(\bm{x},t)=\left[f_{1}(\bm{x},t),\dots,f_{k}(\bm{x},t)\right],$ (1)
where $\bm{f}$ comprises $k$ objective functions, each of which represents a distinct performance measure of an ML model, e.g., validation loss and training cost. Each $\bm{x}$ denotes a $d$-dimensional hyperparameter setting within a compact set $\mathbb{X}\subset\mathbb{R}^{d}$, and $t\in\mathbb{T}=\{1,\dots,t_{max}\}$ the number of epochs used for training. Due to the iterative learning procedure, querying any feasible pair $(\bm{x},t)$ in EMOHPO requires optimizers to sequentially observe the noisy performances $\bm{y}(\bm{x},t^{\prime})=\left[y_{1}(\bm{x},t^{\prime}),\dots,y_{k}(\bm{x},t^{\prime})\right]$ for each epoch $t^{\prime}=1$ to $t$, where $y_{i}(\bm{x},t^{\prime})=f_{i}(\bm{x},t^{\prime})+\varepsilon_{i}$ and $\varepsilon_{i}\sim\mathcal{N}(0,\sigma_{i}^{2})$ with variance $\sigma_{i}^{2}$ for any $i\in\{1,\dots,k\}$. (For simplicity, we adopt the common assumption that the noise terms are i.i.d. across both hyperparameter settings and epochs [Dai et al., 2019, Klein et al., 2022].) In order words, querying $(\bm{x},t)$ results in a sequence of $t$ queries from $(\bm{x},1)$ to $(\bm{x},t)$, which necessitates the design of a more sample-efficient strategy to avoid redundant queries of the same hyperparameter setting with different epoch numbers, especially in decreasing order.
In the context of iterative learning, each objective function $f_{i}(\bm{x},\cdot)$ with fixed hyperparameter setting $\bm{x}$ is generally viewed as a learning curve. However, a learning curve only allows for the analysis of one performance measure at a time. To comprehensively analyze multiple performance measures, we introduce the concept of trajectory for any hyperparameter setting $\bm{x}$ as the collection of all noise-free model performances during the training with $\bm{x}$, i.e.,
$Trj(\bm{x}):=\left\{\bm{f}(\bm{x},t)\right\}_{t=1}^{t_{max}}=\left\{\left[f_{1}(\bm{x},t),\dots f_{k}(\bm{x},t)\right]\right\}_{t=1}^{t_{max}}.$
A trajectory inherently encapsulates the information provided by multiple learning curves. For notational convenience, let $\bm{z}=(\bm{x},t)$ and $\mathbb{Z}=\mathbb{X}\times\mathbb{T}$. We then adopt the standard definitions of Pareto optimality for multi-objective minimization problems.
@@
The Pareto-optimal set of $\mathbb{Z}$, denoted by $Z^{*}$, is composed of $\bm{z}\in\mathbb{Z}$ whose $\bm{f}(\bm{z})$ is not dominated by $\bm{f}(\bm{z}^{\prime})$ of any other $\bm{z^{\prime}}\in\mathbb{Z}$, i.e., $Z^{*}=\{\bm{z}\in\mathbb{Z}\mid\nexists\bm{z}^{\prime}\in\mathbb{Z},\bm{f}(\bm{z}^{\prime})\prec\bm{f}(\bm{z})\}$. The corresponding set of solutions $F^{*}=\{\bm{f}(\bm{z})\mid\bm{z}\in Z^{*}\}$ is referred to as Pareto-optimal front.
As per Definitions 1 and 2, minimizing the EMOHPO in (1) is equivalent to locating the Pareto-optimal set over the entire search space, except that this space is composed of all feasible pairs of hyperparameter settings and training epochs. This aligns with the purpose of this study of finding the trade-offs over trajectories.
Finally, it is important to highlight that, beyond incorporating iterative learning, EMOHPO fundamentally differentiates itself from multi-fidelity MOHPO by including the training epoch as a decision variable. Compared to (1), multi-fidelity MOHPO confines its search to trade-offs among the fully-trained model performances (at the highest fidelity) only and can be expressed as,
$\min_{\bm{x}\in\mathbb{X}}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \bm{f}(\bm{x},t_{max})=\left[f_{1}(\bm{x},t_{max}),\dots,f_{k}(\bm{x},t_{max})\right],$ (2)
where the training epoch $t_{max}$ is a fixed constant. In contrast, EMOHPO broadens the search domain from $\mathbb{X}\times\{t_{max}\}$ in multi-fidelity MOHPO to $\mathbb{X}\times\mathbb{T}$ and thus accounts for all observations on the trajectories. As a consequence, the Pareto-optimal front of EMOHPO is always superior to or at least equivalent to that of multi-fidelity MOHPO because the former additionally captures trade-offs that may emerge during iterative learning (see Figure 1(C)). This broader perspective enables more efficient decision-making in hyperparameter tuning, particularly for scenarios requiring model retraining.
@@ 3 Related Work in Bayesian Optimization
Bayesian Optimization for Iterative Learning: By appropriately characterizing the learning curve, epoch-efficient BO aims to predict the fully-trained model performance based on a partially observed learning curve to avoid ineffective epochs of training. Freeze-Thaw BO [Swersky et al., 2014] introduces GP with an exponential decaying kernel to model the validation loss over time and allows the training to be paused and later resumed under hyperparameter settings that show promise. BOHB [Falkner et al., 2018], a BO extension of HyperBand [Li et al., 2018], allocates training epochs through random sampling and utilizes successive halving to eliminate suboptimal hyperparameter settings. Instead of dynamically allocating computational budget among hyperparameter settings, BO-BOS [Dai et al., 2019] combines BO with Bayesian optimal stopping to early stop the training with a hyperparameter setting predicted to yield poor model performance. Similarly, both BOIL [Nguyen et al., 2020] and BAPI [Belakaria et al., 2023] incorporate strategies for early stopping with a particular focus on considering the learning curve of the training cost. BOIL integrates the cost into the acquisition function to prioritize cost-effective ML training procedures, whereas BAPI imposes a fixed total budget over cost. Unfortunately, there has been no epoch-efficient multi-objective BO.
@@
Multi-Objective Bayesian Optimization: Many MOBO methods have been developed for multi-objective HPO by extending the vanilla BO framework. These methods generally build a surrogate model for each objective to minimize the necessity of actual resource-intensive objective evaluations, and they differ in the implementation of acquisition function. A straightforward approach is to convert a multiple-objective problem into a single-objective problem through techniques like random scalarization, which allows a direct application of standard acquisition functions. For example, ParEGO [Knowles, 2006] and TS-TCH [Paria et al., 2020] respectively apply Expected Improvement (EI) [Jones et al., 1998] and Thompson Sampling (TS) [Thompson, 1933]. However, this approach often encounters limitations in adequately exploring the Pareto-optimal front. Therefore, acquisition functions biased towards the Pareto-optimal front, such as Expected Hypervolume Improvement (EHVI) [Emmerich et al., 2006], have been designed. Despite the effectiveness and popularity of EHVI, its computational intensity presents a significant challenge in the development of BO methods [Hupkens et al., 2015]. More recently, $q$EHVI [Daulton et al., 2020] and $q$NEHVI [Daulton et al., 2021] have extended EHVI for parallel multi-point selection and batch optimization through Monte Carlo (MC) integration [Emmerich et al., 2006], and they have demonstrated notable empirical performance. Alternatives to EHVI can be found in [Hernández-Lobato et al., 2016, Belakaria et al., 2019, Suzuki et al., 2020, Yang et al., 2022, Daulton et al., 2022].
@@
Multi-Fidelity Bayesian Optimization: Multi-fidelity BO has a rich research history [Kandasamy et al., 2016, 2017, Sen et al., 2018, Wu et al., 2020, Fan et al., 2024] and it facilitates the optimization of fully-trained model performance by utilizing its low-fidelity approximations, which can be obtained either by using a partial training dataset or by limiting the number of training epochs. Although multi-fidelity BO shares similarities with epoch-efficient BO, it predetermines the fidelity level before initiating the iterative learning procedure and therefore ignores the observations during the procedure. For example, FABOLAS [Klein et al., 2017] considers the data subset size as the fidelity level and jointly selects the hyperparameter setting and the data subset for model training. BOCA [Kandasamy et al., 2017] expands the discrete fidelity space to continuous for a more general setting. Multi-fidelity MOBO has also been studied in the literature [Belakaria et al., 2020, Schmucker et al., 2021]; however, it is not applicable to solving EMOHPO as defined in (1). This is because multi-fidelity MOBO treats the training epoch merely as an additional degree of freedom, rather than part of the decision variables. Consequently, its objective remains focused on solving multi-fidelity MOHPO as defined in (2).